OpenStudy (amistre64):

A store has 80 parts in its inventory. 30 are from source A, 50 from source B. 20% of A are defective, 8% of B are defective. Whats the probability that exactly 2 from a selection of 5 w/out replacement are defective?

OpenStudy (amistre64):

30(.2) + 50(.08) 6 + 4, 10 defective total P(defect) = 10/80 5C2 = 10, sooo 10(10*9*(78)*(77)*(76))/(80*79*78*77*76) 10(90)/(80*79) whats wrong with tis approach?

OpenStudy (amistre64):

.1424 isnt an option on the practice exam ...

OpenStudy (amistre64):

i dont think it matters where the parts come from after they are in inventory, so picking a part thats defective is 10/80 and 9/79 regardless

OpenStudy (amistre64):

i have 80 30(.2) are bad, and 50(.08) are bad the sources defined the number that are bad from each one. i already know how many are bad regardless of the source since I have all the parts in my possession and i am digging them out of the box that they are all in

OpenStudy (amistre64):

piecing it out shouldnt matter a'a' aaa or a'a' aab or a'a' abb or a'a' bbb or etc .... ive got 80 in my possession, i know that 6 are bad and 4 are bad ... 10 out of 80 are bad whats my error if any?

OpenStudy (amistre64):

i see it, i think 70 are good, not 79

OpenStudy (amistre64):

10(10*9*(70)*(69)*(68))/(80*79*78*77*76) .1024

OpenStudy (rational):



OpenStudy (amistre64):

its a good thing i determined C then lol

OpenStudy (inkyvoyd):

:D

OpenStudy (amistre64):

320am, brain ws collapsing :) 70 and down are good, not 79

OpenStudy (amistre64):

ive got one that is a joint probability function, gonna ask that one next ....and no cheating from nonOS sites lol

OpenStudy (rational):

@amistre64 im getting 0.105 i could be wrong though..

OpenStudy (perl):

I agree with amistre's work You can first calculate the probability of P ( d d d ' d ' d ' ) where d = defective , d ' = non defective, with nonreplacement, and then multiply by 5 choose 2 because P ( d d d ' d ' d ' ) = P( d' d d d' d' ) = ... you can verify this by making a tree. Now P(d d d ' d' d' ) = ( 10 * 9 * 70 * 69 * 68)/ ( 80*79*78 *77*76) , then multiply that by 5 choose 2. and there is confirmation in that link. http://www.actuary.com/actuarial-discussion-forum/archive/index.php/t-10406.html?s=7e3604b14480daf9a5f0ca9eb82ad8a4 I would like to see rational's solution salvaged, that approach could work with some tweaking

OpenStudy (rational):

that link opens a blank page for me :/

OpenStudy (perl):

(10/80) * (10/80) * (70/80) * (70/80) * (70/80) * 5 choose 2 = .1046

OpenStudy (perl):

@rational your solution works "with replacement".